In Hepburn’s Speed Model, the coefficients of vehicles are indicated for C and D. As the chief of operations in your organization, you are responsible for presenting the yearly budget for the semi trucks in your company’s inventory. Since your safety officer is insisting that each of your drivers must maintain an average speed of 55mph, what would be the vehicle operating cost of your company for each semi-truck in cent per mile?
1
Expert's answer
2020-06-15T19:07:32-0400

VOC = a0- a1s + a2s2

In this case s = 55mph

VOC = 38.1 -(0.093 * 55) + (0.00033*55)= 38.1 - (5.115) + (0.99825)= 33.9833